A salesperson's weekly paycheck is 25% less than a second salesperson's paycheck. The two paychecks total $1625. Find the amount of each paycheck. (Round your answers to the nearest cent.)
Found 2 solutions by stanbon, amarjeeth123:
Answer by stanbon(75887)

A salesperson's weekly paycheck is 25% less than a second salesperson's paycheck. The two paychecks total $1625. Find the amount of each paycheck. (Round your answers to the nearest cent.)
-----
Larger paycheck:: x
Smaller paycheck:: 0.75x
----------
Equation:
x + 0.75x = 1625
1.75x = 1625
x = $928.57
0.75x = $696.43
-----------------------
Cheers,
Stan H.
----------------

Answer by amarjeeth123(569)

Let the second person's paycheck be x.
Then the first person's paycheck is (x-25/100x) i.e.3/4x
The sum of the two paychecks is 1625.
x+3/4x=1625
7/4x=1625
x=4*1625/7=928.57
3/4x=696.43
The first person's paycheck is $696 and 43 cents.
The second person's paycheck is $928 and 57 cents.